OpenStudy (some.random.cool.kid):

x y 0 2 1 −1 2 −6 Find the average rate of change for the given function from x = 0 to x = 2.

OpenStudy (some.random.cool.kid):

Find the average rate of change for the given function from x = 0 to x = 2. 2 4 −4 1

OpenStudy (some.random.cool.kid):

@jim_thompson5910 please help with a few of these...ik ur super decent at these types of maths.

jimthompson5910 (jim_thompson5910):

what's the given function?

jimthompson5910 (jim_thompson5910):

oh it's given as a table?

OpenStudy (some.random.cool.kid):

yes

OpenStudy (some.random.cool.kid):

x 0 1 2

OpenStudy (some.random.cool.kid):

with y as

OpenStudy (some.random.cool.kid):

2 -1 -6

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

when x = 0, what is the value of y?

OpenStudy (some.random.cool.kid):

says it 2

OpenStudy (some.random.cool.kid):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so (0,2) is one point on the graph

jimthompson5910 (jim_thompson5910):

when x = 2, what is the value of y?

OpenStudy (some.random.cool.kid):

its -6

OpenStudy (some.random.cool.kid):

there is no graph.. its a drawn out imaginative graph i guess.

OpenStudy (some.random.cool.kid):

uhm u sleep? lol

jimthompson5910 (jim_thompson5910):

you would use the table

jimthompson5910 (jim_thompson5910):

look where x = 2. Then look directly across

OpenStudy (some.random.cool.kid):

i said it was -6 ...

jimthompson5910 (jim_thompson5910):

ok so we have (0,2) and (2,-6) Find the slope of the line through those points

OpenStudy (some.random.cool.kid):

draw it?

jimthompson5910 (jim_thompson5910):

Slope Formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) = (0,2) (x2,y2) = (2,-6)

jimthompson5910 (jim_thompson5910):

\[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{-6-2}{2-0}\] \[\Large m = ???\]

OpenStudy (some.random.cool.kid):

3?

OpenStudy (some.random.cool.kid):

oh wait... -4

jimthompson5910 (jim_thompson5910):

-4 yes

jimthompson5910 (jim_thompson5910):

\[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] \[\Large m = \frac{-6-2}{2-0}\] \[\Large m = \frac{-8}{2}\] \[\Large m = -4\]

OpenStudy (some.random.cool.kid):

ok great thanks

jimthompson5910 (jim_thompson5910):

no problem